OpenStudy (anonymous):

Match the following 1-1 function with its inverse f(x)=5x^3+7

OpenStudy (anonymous):

@Peaches15

OpenStudy (anonymous):

@LynFran let me help her with dis one k

OpenStudy (anonymous):

@enchanted_bubbles help her plzzz

OpenStudy (lynfran):

to get its inverse f(x)=5x^3+7 y=5x^3+7 solve for x... y-7=5x^3 (y-7)/5=x^3 the cube root of (y-7)/5=x now replace x and y we get cube root of (x-7)/5=y'

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

What is the ratio of the number of days in the week that do not begin with the letter "T" to the number of days in the week that do begin with the letter "T"? 7 over 2 2:5 2 to 7 5 over 2